l

Pastry Garden Bakery manufactures three types of baked goods for retail sale. The three types are: wedding cakes, pies and cookies. Currently, Pastry Garden uses a single plant-wide overhead rate to allocate its $1,667 250 of annual manufacturing overhead. Of this amount, $562,000 is associated with wedding cakes, $492,000 is associated with pies, and $612 750 is associa with cookies. Pastry Garden is currently using a total of 22,150 labor hours: 7.500 for wedding cakes, 8.200 for pies, and 6,450 fo cookies. Pastry Garden uses labor hours to allocate overhead costs What is the plant wide overhead rate? What is the departmental overhead rate for wedding cakes? What is the departmental overhead rate for pies? What is the departmental overhead rate for cookies? Which product line is undercosted by using the plant wide manufacturing By how much per machine hour is it undercosted?
Answer #1
The plant wide overhead rate = Total annual Manufacturing overhead / Total labor hours
The plant wide overhead rate = $1667250 / 22150 LHs = $75.27 per LH
Departmental overhead rate for wedding cakes = Manuf O/H associated with wedding cakes / LH for wedding cakes
Departmental overhead rate for wedding cakes = 562000/ 7500 = $74.93 per Wedding cakes LH
Departmental overhead rate for pies = 492000 / 8200 = $60 per pies LH
Departmental overhead rate for cookies = 612750 / 6450 = $95 per cookies LH
Cookies product line is being undercosted if plant wide manuf. Overhead rate is used because
cookies departmental overhead rate actually is $95 per cookies LH.
Cookies product lien has been undercosted by = $95 - 75.27 = $19.73 per machine hour